### v2

# A multi-objective mixed integer linear programming model for supply chain planning of 3D printing

Abstract
3D printing is considered the future of production systems and one of the physical elements of the Fourth Industrial Revolution. 3D printing will significantly impact the product lifecycle, considering cost, energy consumption, and carbon dioxide emissions, leading to the creation of sustainable production systems. Given the importance of these production systems and their effects on the quality of life for future generations, it is expected that 3D printing will soon become one of the global industry’s fundamental needs. Although three decades have passed since the emergence of 3D printers, there has not yet been much research on production planning and mass production using these devices. Therefore, we aimed to identify the existing gaps in the planning of 3D printers and to propose a model for planning and scheduling these devices. In this research, several parts with different heights, areas, and volumes have been considered for allocation on identical 3D printers for various tasks. To solve this problem, a multi-objective mixed integer linear programming model has been proposed to minimize the earliness and tardiness of parts production, considering their order delivery times, and maximizing machine utilization. Additionally, a method has been proposed for the placement of parts in 3D printers, leading to the selection of the best edge as the height. Using a numerical example, we have plotted the Pareto curve obtained from solving the model using the epsilon constraint method for several parts and analyzed the impact of the method for selecting the best edge as the height, with and without considering it. Additionally, a comprehensive sensitivity and scenario analysis has been conducted to validate the results.
keywords:
Operations research, supply chain planning, scheduling, job allocation, part orientation optimization, 3D printing
[…]
organization=Department of Integrated Systems Engineering, addressline=The Ohio State University, city=Columbus, postcode=43210, state=OH, country=USA
1 Introduction
The increase in population and the growing need for natural resources such as water, energy, food, etc., to meet the needs of the rising population, have prompted the current generation to seek solutions for fulfilling the needs of future generations. Many of the current methods used to meet human needs are not sustainable and have adverse effects on the environment, economy, human health, etc. Therefore, many scientists and researchers are looking for ways to meet human needs that are renewable and have minimal long-term negative impacts on the environment and the life cycle. 3D printing, as one of the newest production systems in recent decades, has garnered much attention from researchers and industry professionals and is rapidly advancing. This new technology, utilizing minimal raw materials and energy, is considered the future of mass production systems. It can meet many human needs in industrial, medical, food, aerospace, and other fields sustainably and cost-effectively. The significance of this topic in these areas has led to extensive research for the commercialization and industrialization of this technology. Since production planning in supply chains is one of the significant issues for significantly reducing costs, production time, and so on, the production planning of 3D printers is crucial in the implementation of this technology in the industry. Many old-fashioned industrial companies can revamp their old production systems and benefit from the numerous advantages of 3D printing technology, such as high flexibility in product manufacturing, low cost, high accuracy, and so forth, by investing in this technology. 3D printers, with only digital information of the final products, can produce consumer goods in a short time and small quantities. This aspect can be beneficial for the companies in several ways. For example, products are manufactured exactly according to the specifications ordered by customers, which is a significant advantage in competitive markets. Additionally, the proximity of production centers using additive technology to final consumers reduces transportation costs. Optimal energy consumption, raw materials, and the need for fewer employees lower the final product price, which is a major competitive advantage. The significant question raised in this field is: How can we maximize the benefits of 3D printing technology? given the importance of production planning in manufacturing systems, how can we implement this effectively in these systems? It is important to note that 3D printers have only begun their industrialization process in recent years, as they still do not have high speeds for mass production and currently have fewer advantages compared to mass production systems. However, with the significant advancements in these machines over the past decade, it is expected that many organizations will replace conventional production systems with this system in the coming years. One of the best production approaches is for manufacturers to deliver customer orders exactly on time and essentially maintain zero inventory, which reduces the total cost of goods. The costs of purchasing, maintaining, and operating 3D printers are high because they have not yet reached the stage of mass production. Therefore, planning that optimizes the use of these devices and ensures timely delivery of orders will reduce production, maintenance, and lost profit costs. The main goal of this paper is to find such a plan for 3D printers. To address this problem, we proposed a multi-objective mixed integer linear programming (MOMILP) model. To the best of our knowledge, there exists no well-known work integrating the just-in-time concept of part manufacturing combined with part orientation optimization and maximization of 3D printer utilization. In particular, the orientation of parts to be produced by 3D printers can cause delays in order satisfaction and machine utilization rate. Our model includes the part orientation optimization and also maximization of 3D printer utilization. Indeed, the goal of our model is to maximize the utilization of machines and minimize the earliness or tardiness of parts manufacturing including part orientation optimization. Finally, we conducted several experiments on randomly generated data to validate our model. Several managerial insights have been concluded from the results, i.e., part orientation optimization can significantly reduce the earliness and tardiness of the production of the part given the available surface area of the printers. Similarity and discrepancies in dimensions and due dates of parts in a part are important in reducing the sum of earliness and tardiness of parts productions highlighting the significance of job allocation and sequencing.
The outline of this paper is as follows: In section 2, we provide some background on the concepts due to the complexity of the problem and provide some definitions. A comprehensive literature review has been gathered in section 3. Section 4 defines the problem, section 5 elaborates on the MOMINLP model, and section 6 provides numerical experiments and sensitivity analysis. Finally, section 7 concludes and proposes future avenues.
2 Background
Before discussing 3D printing, it is necessary to choose a term to refer to it. Many terms have been used to describe 3D printing due to the rapid advancement of 3D printing technology. Over time, new terms have emerged to describe its new applications.
Rapid prototyping: The term ”rapid prototyping” refers to the swift creation of a component before its final market release. Because layered manufacturing was originally employed to produce prototypes, it became associated with the term rapid prototyping.
Rapid manufacturing: After a while, 3D printing technology began to be used for the direct production of final products. At this stage, the term rapid prototyping could no longer describe this new capability. The new term rapid manufacturing better described this capability.
Rapid tooling: Subsequently, this technology was used in traditional manufacturing processes. Therefore, the term rapid tooling was used to describe 3D printing processes [1].
In early efforts to name additive technology and its applications, the term layered manufacturing was chosen as a name for this technology used for producing 3D products as assembly parts or final products. Unlike subtractive or traditional manufacturing, additive systems bond liquids, powders, or metal sheets to create parts. These parts may be structurally complex and difficult to produce using other technologies. This definition highlights the different features of 3D printing compared to traditional manufacturing systems [2].
2.1 History
3D printing technology began to grow rapidly in the late 1990s. However, this technology entered a competitive phase with traditional manufacturing after a two-decade delay and witnessed significant technical and entrepreneurial changes. Initially, this technology was used in industrial design markets before products reached mass production. This technology allowed manufacturers to produce parts in limited quantities much faster. With 3D printing, initial prototyping, which previously required several months or weeks with traditional manufacturing methods due to the need for molds and lengthy production lines, could be completed within hours or days. Moreover, 3D printers enabled designers to create parts with greater structural and geometric complexity. As a result of these advancements, improvements in quality, cost, and time have made 3D printing a leading method for prototyping [3].
2.2 The Fourth Industrial Revolution
The world has experienced three major industrial revolutions. The first one, in the 18th century, was driven by the invention of the steam engine. The second revolution, at the end of the 19th century, was characterized by the development of electricity and mass production techniques. The third revolution began in the 1970s, marked by advances in communication and telecommunication systems. The concept of the Fourth Industrial Revolution, or Industry 4.0, was coined at the Hannover Fair in 2011. Both public and private organizations have been working to expand this new model of production [3]. Four major factors have been identified as driving forces behind Industry 4.0 [4, 5]:
  1. 1.
    Significant increases in data volume, computing power, and communication capabilities.
  2. 2.
    The rise of advanced analytics and business intelligence.
  3. 3.
    New forms of interaction between humans and machines.
  4. 4.
    The improved transfer of digital data to the physical world.
Research by [3, 6] has highlighted six key factors expected to influence the future economy, industry, and technology:
  1. 1.
    Heavy reliance on the internet.
  2. 2.
    Increased communication, data processing, and storage needs.
  3. 3.
    Expansion of material networks.
  4. 4.
    Advances in artificial intelligence and information sciences.
  5. 5.
    Growth of the shared economy.
  6. 6.
    Continued digitization.
According to predictions by [3, 6], significant events related to 3D printing were anticipated to occur between 2018 and 2024. Their study also projected several milestones for 3D printing by 2025:
  • 1.
    The production of the first car entirely manufactured using 3D printing (84% consensus).
  • 2.
    5% of consumer products being made with 3D printing technology (84% consensus).
  • 3.
    The creation of the first intestinal transplant using 3D printing methods (76.54% consensus).
[3, 6] reported that over two decades of 3D printer commercialization led to a cumulative annual growth rate of 25.4% in global revenue for 3D printing products in 2013. The growth rate rose by 27.4% from 2010 to 2012, achieving \$2.2 million in 2012. The sales of 3D printers grew by 19.3% in 2012, and the purchases of personal 3D printers saw a 46.3% rise in the same year, reflecting the growing popularity of 3D printers among both retailers and consumers. [7] conducted another study examining the factors for using 3D printers. They did a global survey with 330 participants, who were workers of organizations with over 100 staff, either utilizing 3D printers or intend to implement them. They showed that half of these organizations were using this technology in the early stages of production to improve their manufacturing processes. Interestingly, 6% of these individuals believed that the high cost of this technology was a barrier to its implementation, although companies currently using this technology believed that it had many benefits and that if the costs of implementation were lower and the technology improved, many companies would adopt this technology. Currently, over 67% of organizations are using this technology, and this percentage is expected to increase significantly in the future.
2.3 Applications
In its evolution, 3D printing has progressed through three distinct phases. Initially, product designers utilized this technology to create prototypes efficiently and affordably, ensuring rapid production and enhanced security. Subsequently, 3D printing advanced to encompass final product manufacturing, significantly reducing supply chain length, and delivery times, and fostering innovation essential for competitive advantage. Looking forward, there is anticipation that consumers will increasingly adopt 3D printing for personal product creation in the near future [3].
2.3.1 Industry Applications
The concept of Industry 4.0 sparked efforts by both governments and private enterprises to develop this new industrial paradigm [3]. Sustainability, a key topic today, involves achieving economic success, environmental conservation, and social equity simultaneously [8]. Sustainable manufacturing aims to use resources efficiently while reducing emissions and energy consumption. As we face future challenges like fuel shortages, volatile prices, limited raw materials, climate change, pollution, and space limitations, the importance of 3D printing technology becomes evident. This technology offers promising solutions to these pressing issues [9].
2.3.2 Medical Applications
Doctors at the Women and Children’s Hospital are using medical models produced by 3D printers to help children and patients with congenital heart diseases and complex conditions. Currently, 3D-printed medical models have become an integral part of treatment plans worldwide and are even used in hospitals for treating head and bone injuries. In recent years, hospitals relied on 2D scans for diagnosing and treating heart diseases, using MRI scans for heart diagnosis like most doctors. However, they now have basic structures for medical models. They can simply use data through 3D modeling software and 3D printers to construct the models within two hours. Researchers at MIT pioneered this process, and now doctors worldwide can have a medical model as part of their standard diagnostic package. They have proven the value of this method, which has been implemented in hospitals for women and children. Doctors have found that the benefits of medical 3D printing go beyond simple protection of the patient’s damaged heart. Doctors can cut the model and plan surgery based on the specific defects of each individual. This method is effectively applied in real surgeries, providing doctors with an opportunity to solve potential problems on 3D-printed medical models. This opportunity for surgical planning has led to advancements, meaning doctors spend less time in the operating room, patients recover faster, and significant cost savings are realized [10].
2.3.3 Aerospace Applications
A 3D printer has been currently launched in the International Space Station (ISS) for onboard experiments used on the exterior of the ISS. This 3D printer, produced by NASA and used in space, can work with polymer filaments and perform rapid prototyping in small sizes for producing spare parts for satellites. In recent months, NASA announced the construction of a 3D printer named AMF, which is used in space. AMF can also produce tools and provide maintenance processes, especially offering more space manufacturing flexibility [3, 11].
2.4 Types of 3D Printers
3D printers can be divided into 18 categories, which can be grouped into three main categories [12]:
  1. 1.
    FDM (Fused Deposition Modeling): This process involves injecting thermoplastic materials in filament form for 3D printing, which reached a commercial stage in 1991.
  2. 2.
    SGC (Solid Ground Curing): A UV-sensitive polymer liquid shapes all layers in one pass of UV light over a glass plate with an electrostatically shaped powder coating.
  3. 3.
    SLS (Selective Laser Sintering): Introduced in 1992, this technology uses lasers to fuse powdered materials.
Since 1994, various improved models have been created by inventors.
2.5 Operations research concepts and definitions
In this section, we define general concepts and fundamental terms used throughout the project.
2.5.1 Production Planning
Production planning involves the strategic management of work and workloads within the production process, encompassing setting, controlling, and optimizing these activities. Factories employ both forward and backward planning techniques to allocate resources, plan human resources, schedule production tasks, and procure materials. Batch production planning specifically pertains to scheduling batch production processes. While scheduling is crucial for continuous traditional processes like refining [13], it is particularly critical for batch processes such as ”active pharmaceutical ingredients, biotechnology processes, and specialized chemical processes” [14]. Batch production planning shares methodologies with capacitated production planning, which addresses various production-related challenges [15]. Mathematical scheduling methods approach production planning as an optimization problem, aiming to minimize or maximize specific objectives under a set of constraints typically represented as equations. These objectives and constraints may involve binary variables, integers, and non-linear relationships. Constraint programming provides a systematic approach to solving such problems by treating them as a set of constraints, striving for efficient solutions in a timely manner [16]. For instance, [17] explored revenue management focusing on intricate pricing dynamics. They utilized a constrained MILP model to enhance revenue optimization, using product prices as decision variables while integrating capacity and consumer behavior within the model’s constraints.
2.5.2 Job sizing and Scheduling
Job sizing and scheduling involve optimizing the assignment of jobs to resources at specific times [18]. In this context, the objective is to schedule $n$ jobs of varying sizes across $m$ machines to minimize total manufacturing time. Today, this problem is considered dynamic scheduling, requiring algorithmic decisions based on available information before each job’s scheduling [15, 19].
Various types of job scheduling problems can be categorized as follows [15, 18, 19]:
  1. 1.
    Machines may be dependent, independent, or identical.
  2. 2.
    Machines may introduce specific job sequencing or perform jobs without gaps.
  3. 3.
    Machines may be sequenced based on scheduling sequences.
  4. 4.
    Objectives may include minimizing total time, minimizing job tardiness, maximizing job delay, and more.
  5. 5.
    Jobs can have constraints, such as job $i$ needing to finish before job $j$ starts.
  6. 6.
    Shared constraints by jobs and machines restrict certain jobs to specific machines.
  7. 7.
    A set of jobs may be associated with different sets of machines.
  8. 8.
    Exact or probabilistic processing times.
This paper addresses a combined problem involving cases 1, 2, 4, 5, and 8.
2.5.3 Bin-packing problem
The bin-packing problem deals with placing parts of different sizes in a limited number of machines, using the minimum possible number of machines. Such problems are characterized by non-deterministic polynomial-time hardness [20]. This decision-making problem (deciding on parts to be placed in a specific number of machines) is one of the most difficult problems in this category and belongs to the NP-complete problem category [21]. Non-deterministic polynomial-time hardness in computational complexity theory is among the most difficult polynomial-time problems. A simple example of non-linear polynomial-time hardness problems is the subset sum problem [22]. Various types of these problems are discussed, such as linear packaging, weight-based packaging, cost-based packaging, and so on, which have applications in loading containers, trucks, and so on with weight constraints, etc.
2.5.4 Multi-objective Decision Making
Multi-objective optimization problems involve multiple objective functions [23]. Mathematically, a multi-objective optimization model is formulated as follows [24, 25]:
$\min(f_{1}(x),f_{2}(x),\ldots,f_{k}(x))$
$subject\ to\ x\in X$
where $k\geq 2$ is the number of objectives in the model and $X$ is the feasible set of decision vectors. Additionally, the objective function vector is defined as follows:
$f:X\to\mathbb{R}^{k},\quad f(x)=(f_{1}(x),\ldots,f_{k}(x))^{T}$
If we aim to maximize certain objective functions, it implies minimizing the negative of those objective functions. An element $x^{*}\in X$ is termed a feasible solution. The vector $z^{*}=f(x^{*})\in\mathbb{R}^{k}$ for a feasible solution $x^{*}\in X$ is referred to as the outcome or objective vector. In multi-objective optimization, achieving a solution that minimizes all objective functions at the same time is generally not feasible. Instead, we focus on Pareto optimal solutions. Pareto solutions are solutions where improving some objective functions comes at the cost of others, and it’s impossible to improve all objectives simultaneously. In mathematical terms, a feasible solution $x^{1}\in X$ is said to dominate a solution $x^{2}\in X$ if:
$f_{i}(x^{1})\leq f_{i}(x^{2})\quad\forall i$
$f_{j}(x^{1})<f_{j}(x^{2})\quad\text{for at least one index}$
A solution $x^{*}\in X$ is Pareto optimal if there is no other solution that dominates it. The collection of all Pareto optimal solutions forms the Pareto front. In a multi-objective optimization model, the Pareto front is defined by the objective vectors corresponding to the worst solutions. , defined as follows:
$Z_{i}^{nad}=\sup_{x\in X}f_{i}(x)$ $\forall i\in\{1,..,k\}$
The ideal vector is defined as follows:
$Z_{i}^{ideal}=\inf_{x\in X}f_{i}(x)$ $\forall i\in\{1,..,k\}$
Put differently, the components of the ideal and worst objective vectors define the upper and lower bounds of the Pareto optimal solutions for an objective function. In practical terms, estimating the worst objective vector is necessary because the exact Pareto optimal set is often unknown. For this reason, the utopian objective vector $z^{\text{utopian}}$ is defined as follows:
$z_{i}^{\text{utopian}}=z_{i}^{\text{ideal}}-\epsilon\quad\forall i=1,\ldots,k$
where $\epsilon>0$ is a small constant defined for computational reasons. Several methods have been suggested for optimizing multi-objective problems, including Lexicographic Programming, Lexicographic Minimization Approach, Lagrangian relaxation, and the $\epsilon$-constraint method, among others. This paper specifically employs the $\epsilon$-constraint method. In the $\epsilon$-constraint method, one objective function is chosen for optimization, while the remaining objective functions are transformed into constraints. These constraints are bounded by upper limits, which can be interpreted as values of Reference Levels (RLs). The essence of the $\epsilon$-constraint problem is outlined as follows [26]:
$\min f_{j}(x)$
$s.t.$
$f_{i}(x)\leq\epsilon_{i}$ $\forall i=1,2,...,k\ \&\ j\neq i\text{ (}\epsilon_{i}=RL\text{)}$
2.5.5 Part Orientation
Part orientation during 3D printing affects the accuracy, resistance, surface quality of the part, and manufacturing speed. For instance, Consider a cylinder with a hole in the middle that is to be printed vertically with a 3D printer. The 3D printer creates the component by placing a series of concentric circles on top of each other. This process creates a final cylinder with a smooth outer surface. If the orientation of this cylinder is in the direction of the horizon (y-axis or x), the part is made up of several polygons that are made layer by layer. The surface of the cylinder, which interfaces with the construction tray, is smoothed during manufacturing. The orientation of the part significantly influences manufacturing time; for instance, producing a cylinder horizontally is more time-consuming and costly than vertically. Layer-by-layer construction in 3D printing results in parts having greater strength in the x-y direction rather than in the z-direction. This orientation disparity refers to how the part is positioned during printing. For example, the likelihood of breakage along the z-axis can be up to five times higher than along the x-y axes. Generally, parts made with 3D printers exhibit greater strength along one axis compared to others [3, 27].
3 Literature Review
Many steps are involved in transforming raw materials into final products within a production chain. Over recent decades, the emergence of 3D printing technology has revolutionized these steps in final product production, including design, planning, manufacturing, and sales. 3D printing enables unique products to reach customers in both small and large markets without the need for inventory or high production costs. Production planning is a framework that establishes production objectives, identifies necessary production resources, and formulates plans to achieve production goals on time. It anticipates future risks and prepares various scenarios to minimize waste. With the advent of 3D printing, production planning has transformed. One significant change is the reduction in the number of suppliers, as the amount of consumables in 3D printing is lower compared to traditional industrial production. In layer-by-layer production, parts produced are either final products or components of another assembly, ready for immediate use, whereas traditional production involves multiple stages to produce a single piece [28]. Quantitative research has focused on production planning and scheduling for 3D printing systems to optimize machine utilization. [29] proposed a method using MILP models to meet customer demand within delivery time windows, thereby reducing delays. Although their study does not address part orientation optimization, their primary objective is to minimize delivery delays. They also aim to minimize the time differential in delivering orders within a batch. [30] presented a mathematical model for scheduling SLM machines, an NP-hard problem aiming to minimize lateness and associated costs. To tackle the complexity, their paper proposes metaheuristic algorithms, specifically, Genetic Algorithms adapted for matrix handling. The capabilities of these algorithms are validated through a case study that includes a production plan with 30 parts. [31] tackled the challenge of printing uniquely configured parts across multiple machines with specific deadlines. The objective function was to minimize total time while adhering to all deadlines, utilizing a mixture of constrained bin packing, two-dimensional nesting, and job shop scheduling, framed through constraints and graph theory to create a formal model. [32] introduced a decision support model employing multi-objective optimization for managing batches of parts across several machines. The focus was on balancing operating costs, printer load, total tardiness, and the number of defective parts in the Fused Deposition Modeling (FDM) process. They validated the model by a case study involving automotive parts, and a trade-off analysis was conducted to examine conflicting objectives. [33] explored the maintenance and repair of 3D printers, noting that in the event of a sudden failure, the printer must restart the production process from the beginning after repairs. They proposed a stochastic preemptive-repeat scheduling model, establishing measures to minimize expected completion and flow times and solving an optimization problem to identify optimal job sizes for printing, specifically addressing dental 3D printing manufacturing challenges. [34] explored scheduling challenges for both single and multiple 3D printers by developing mathematical optimization models. They used MILP to assign parts to jobs on 3D printers, aiming to minimize makespan. Their approach considered various machine configurations such as single machines, parallel identical machines, and parallel non-identical machines. The models were solved using the CPLEX solver, providing optimal solutions within specific time limits (1800 and 2400 seconds). The study focused on scheduling issues involving parallel identical and non-identical Additive Manufacturing (AM) machines. [35] investigated the optimal allocation of parts to machines with diverse specifications, including production time, material cost per volume, processing time per unit height, setup time, maximum supported area, and height. They introduced a mathematical model implemented in CPLEX and devised heuristic methods (’best-fit’ and ’adapted best-fit’ rules) to achieve efficient solutions. The heuristics were explained step by step with examples, and numerical tests demonstrated their effectiveness in reducing 3D printing processing costs through efficient planning. [36] introduced a strategy for managing dynamic order acceptance and scheduling in Powder Bed Fusion (PBF) systems for on-demand production. They aimed to maximize average profit per unit time by integrating bin packing, batch processing, dynamic scheduling, and decision-making into a metaheuristic approach. Experimental studies validated the profitability of their method for complex, NP-hard scheduling problems. [37] developed methods to optimize the arrangement of multiple parts in 3D printing. They employed a ”3D printing feature-based orientation optimization” to improve production quality by adjusting part orientations. Their ”parallel nesting” algorithm used part projection profiles to maximize efficiency, reducing overall build time and costs. [38] addressed part grouping and scheduling challenges in 3D printing with non-identical parallel machines. They proposed grouping parts by material and optimizing build area usage within machine capacity constraints. Their approach aimed to minimize makespan using a MILP model and evaluated three metaheuristic algorithms for efficiency and scalability. [39] explored scheduling problems in metal 3D printing using multiple non-identical Selective Laser Melting (SLM) machines, which varied in dimensions, speed, and cost. The parts also differed in width, length, height, volume, release date, and due date. The aim was to minimize total tardiness while ensuring that parts were allocated to build platforms without overlapping. They proposed a genetic algorithm to efficiently address this complex scheduling issue. [40] tackled a novel challenge in 3D printing scheduling involving unrelated parallel machines. They considered practical constraints such as two-dimensional packing and varying part release times, compounded by multiple part orientation options that influence processing duration. Their approach utilized a MILP model to minimize makespan, recognizing the problem’s NP-hard complexity. For larger-scale instances, they recommended an Adaptive Large Neighborhood Search (ALNS) algorithm. [41] investigated the impact of build orientation policies on production time in 3D printing, focusing on mass customization models. They evaluated two primary strategies: the Laying Policy (LP), which reduces individual build times by minimizing part height but may increase overall completion times with more parts, and the Standing Policy (SP), which reduces job numbers by lowering base plane projections, although each job takes longer due to increased part height. Numerical experiments using Stereolithography (SLA) indicated that SP could effectively reduce completion time for approximately 40 parts. [42] addressed the scheduling of unrelated parallel batch processing machines in Additive Manufacturing (AM), integrating part orientation selection to minimize makespan. They proposed a MILP model and developed a simulated annealing algorithm with packing strategies, highlighting the benefits of considering multiple part orientations in machine scheduling compared to fixed orientation scenarios. [43] conducted a comprehensive analysis of nesting and scheduling challenges in 3D printing, using bibliometric and systematic review methods to map the field’s evolution. They focused on scheduling and allocating parts to 3D printers to minimize the earliness and tardiness of deliveries while maximizing machine utilization through optimal part orientations. Their novel Multi-Objective Mixed Integer Linear Programming (MOMILP) model was validated through extensive numerical experiments and sensitivity analyses. Based on this recent review paper and the reviewed literature above, the problem addressed in this paper is unique. We address scheduling and allocating parts to 3D printers, focusing on two objectives: minimizing earliness 10 and tardiness of part delivery and maximizing machine utilization by selecting the best part height, that is selecting the orientation of the part in the 3D printers. To this end, we propose a novel MOMILP model and analyze and validate the model through several numerical experiments and sensitivity analysis.
4 Problem Definition
We intend to allocate several parts with specific delivery times to several identical 3D printers so that the parts are produced with minimal earliness and tardiness in delivery time while maximizing the utilization of printers. These two objectives indirectly lead to cost reduction, including maintenance costs, scarcity costs, and 3D printer utilization costs. The 3D printer is of the powder bed type and can only produce one item at a time during production. The capacity of each printer depends on dimensions such as surface area and height, such that if the dimensions of a part are incompatible with the 3D printer’s capabilities, it cannot be produced. Each part is assigned jobs based on its dimensions and delivery due time.
Figure 1: Taken from [34], Allocation of parts to 3D printers and sequencing of jobs
Indeed, as demonstrated by Figure 1, we assign the parts to jobs and printers in such a way that customer orders are delivered just on time (minimizing the earliness or tardiness of part delivery), utilizing the printers efficiently given the part orientation optimization and due dates constraints. We call the utilization of a 3D printer efficient if each job utilizes the maximum possible capacity of a 3D printer, i.e., the maximum available surface area in the 3D printer. Furthermore, to simplify the problem, we assume that the required materials for manufacturing parts are of the same type, and the parts are enclosed in cubic shapes to facilitate the calculation of their surface area. The preparation time for each job is combined with the manufacturing time of the parts. The mathematical model can consider one of the three dimensions of the cubic enclosure as the part’s height, depending on how the part is positioned in the 3D printer. Since 3D printers create parts layer by layer, the 3D printer operates at two speeds: volumetric speed and layering speed. Therefore, selecting the appropriate orientation for parts in the 3D printer and consequently choosing the correct part height for timely delivery is highly effective. Because the maximum height in a task determines the layering time, the model must select both an appropriate combination of parts based on their delivery time and an optimal height for them when arranging parts. For example, Figure 2 illustrates a case in which parts $P_{2}$ and $P_{4}$ determine the completion of the job while the other parts have been completed earlier. This can cause delays and inefficiency in the utilization of the printers. Figure 3 demonstrates a case in which this problem has been tackled by the correct combination of parts in a job. Note that the printer completes all parts layer by layer.
Figure 2: Taken from [44], combination 1. Parts $P_{2}$ and $P_{4}$ determine the completion of the job while the other parts have been completed earlier. This can cause delays and inefficiency in the utilization of the printers
Figure 3: Taken from [44], combination 2. This figure demonstrates a case in which the problem mentioned in Figure 2 has been tackled by the correct combination of parts in a job.
It is also assumed that each part, in all three possible orientations in the printer (cuboid faces), has the best quality and can only be placed in three orientations, with no rotation around the vertical axis, see Figure 4 as a way of illustration.
Figure 4: Taken from [44], part orientation
To solve this problem, a bi-objective mathematical model has been proposed. In the following section, we explain the model precisely.
5 Mathematical Modeling
The developed mathematical model is a multi-objective mixed integer non-linear programming (MOMINLP) model. After the introduction of the sets, parameters, variables, and the MONILP model, we will linearize the model so it can be solved efficiently with known methods.
5.1 Sets
Set of parts $I\ \text{indexed by $i$}$
Set of jobs $J\ \text{indexed by $j$}$
Set of machines $M\ \text{indexed by $m$}$
5.2 Input parameters
Machine area $A_{m}$
Machine height $H^{m}$
Initial height of each part $h_{i}$
Initial width of each part $w_{i}$
Initial length of each part $l_{i}$
Earliness penalty coefficient $Ce$
Tardiness penalty coefficient $Ct$
Due date of each order $d_{i}$
Time to move one millimeter vertically $Ht$
Production time of one cubic millimeter of a part $Vt$
Big M notion $M^{\prime}$
5.3 Decision Variables
All the variables are greater than or equal to zero:
Surface area occupied by part $i$ in a machine $A^{{}^{\prime}}_{i}$
Completion time of job $j$ on machine $m$ $C_{jm}$
Completion time of part $i$ $C^{{}^{\prime}}_{i}$
Earliness duration of part $i$ $E_{i}$
Tardiness duration of part $i$ $T_{i}$
Optimal height of part $i$ $H^{{}^{\prime}}_{i}$
Maximum height in job $j$ on machine $m$ $H^{"}_{jm}$
Processing time of job $j$ on machine $m$ $P_{jm}$
Start time of job $j$ on machine $m$ $s_{jm}$
Binary variable for determining the height of part $i$ $b_{i}$
Binary variable for determining the height of part $i$ $f_{i}$
Linearization variable $lp_{ij}$
Linearization variable $lp^{{}^{\prime}}_{ijm}$
Binary decision variable for assigning part $i$ to job $j$ on machine $m$ $x_{ijm}$
Binary decision variable for assigning job $j$ to machine $m$ $y_{jm}$
First objective function $z$
Second objective function $zz$
5.4 MOMINLP Model
In the following bi-objective model, we aim to minimize the total tardiness and earliness of parts production and maximize the utilization of machines in each job by the optimal allocation of parts to jobs, jobs to machines, the optimal orientation of parts in each job, and occupying the maximum available area in each machine respecting the constraints.
$z=\min\sum_{i}T_{i}Ct+E_{i}Ce$ (1)
$zz=\min\sum_{j,m}A_{m}y_{jm}-\sum_{i,j,m}A^{{}^{\prime}}_{i}x_{ijm}$ (2)
$\sum_{j,m}x_{ijm}=1$ $\forall i$ (3)
$\sum_{i}x_{ijm}\leq y_{jm}M^{{}^{\prime}}$ $\forall j,m$ (4)
$b_{i}+f_{i}\leq 1$ $\forall i$ (5)
$H^{{}^{\prime}}_{i}=(1-b_{i}-f_{i})h_{i}+b_{i}l_{i}+f_{i}w_{i}$ $\forall i$ (6)
$A^{{}^{\prime}}_{i}=b_{i}w_{i}h_{i}+f_{i}h_{i}l_{i}+(1-b_{i}-f_{i})l_{i}w_{i}$ $\forall i$ (7)
$H^{{}^{\prime}}_{i}\leq H^{m}$ $\forall i$ (8)
$\sum_{i}(A^{{}^{\prime}}_{i}\times x_{ijm})\leq A$ $\forall j,m$ (9)
$H^{{}^{\prime}}_{i}-(1-x_{ijm})M^{{}^{\prime}}\leq H^{"}_{jm}$ $\forall i,j,m$ (10)
$P_{jm}=HtH^{"}_{jm}+Vt\sum_{i}l_{i}w_{i}h_{i}x_{ijm}$ $\forall j,m$ (11)
$C_{jm}+P_{j+1,m}\leq C_{j+1,m}$ $\forall j,m$ (12)
$P_{1,m}\leq C_{1,m}$ $\forall m$ (13)
$C^{{}^{\prime}}_{i}=\sum_{j,m}x_{ijm}C_{jm}$ $\forall i$ (14)
$C^{{}^{\prime}}_{i}-d_{i}\leq T_{i}$ $\forall i$ (15)
$d_{i}-C^{{}^{\prime}}_{i}\leq E_{i}$ $\forall i$ (16)
$\sum_{i}x_{i,j+1,m}\leq M^{{}^{\prime}}\sum_{i}x_{ijm}$ $\forall j,m$ (17)
The primary objective (1) aims to minimize the costs associated with both earliness and tardiness in part production. Early delivery incurs storage costs, while tardiness can lead to customer dissatisfaction or penalties, with their impacts adjusted by respective cost parameters. The secondary objective function (2) focuses on maximizing machine space utilization by optimizing part allocation to jobs and jobs to machines, considering optimal part orientations and respecting all constraints. Constraints (3) ensure each part is assigned to exactly one job on one machine. Part assignment to a job on a machine is contingent upon the job being allocated to that machine, governed by constraints (5). Constraints (5) to (9) dictate part orientation and placement within the printer, ensuring only one base side of the part is chosen, determining its vertical dimension, calculating base area, limiting part height within printer constraints, and enforcing printer capacity limits. Constraint (10) selects the maximum height among parts within a job, while (11) computes job processing times on machines without considering setup times. Printing time vertically and volumetrically is computed accordingly. Constraints (12) ensure job completion times on a machine follow proper sequencing, with subsequent job completion times greater than or equal to previous ones plus respective processing times. Initial job completion times on each machine are established by (13). Constraint (14) links part completion times to their respective job completion times. Tardiness (15) and earliness (16) for each part are computed, ensuring non-negative earliness. Finally, (17) enforces proper job sequencing on the same machine.
5.5 Linearization
Note that in the above non-linear model, the second objective function and constraints 14 are the roots of non-linearity as two variables have been multiplied together. To linearize the second objective function we introduce a new variable $lp_{ijm}=A^{{}^{\prime}}_{i}x_{ijm}$. Thus, the second objective function will be:
$zz$ $=\sum_{j,m}A^{{}^{\prime}}_{m}y_{jm}-\sum_{i,j,m}lp_{ijm}$ (18)
which is linear now. To engage these variables in the model, we need to introduce them to the constraints. To do so, firstly we need to replace 9 with the following constraints:
$\sum_{i}lp_{ijm}\leq A$ $\forall j,m$ (19)
Additionally, the following set of constraints will ensure that $lp_{ijm}=A^{{}^{\prime}}_{i}$ if $x_{ijm}=1$ and 00 otherwise in a linear fashion:
$lp_{ijm}\leq M^{{}^{\prime}}x_{ijm}$ $\forall i,j,m$ (20)
$lp_{ijm}\leq A^{{}^{\prime}}_{i}$ $\forall i,j,m$ (21)
$A^{{}^{\prime}}_{i}-(1-x_{ijm})M^{{}^{\prime}}\leq lp_{ijm}$ $\forall i,j,m$ (22)
To linearize the constraints 14, we introduce the positive variables $lp^{{}^{\prime}}_{ijm}=x_{ijm}C_{jm}$. As above, to engage these variables in the model, we need to introduce them to the constraints. To do so, firstly we need to replace 14 with the following constraints:
$C^{{}^{\prime}}_{i}=\sum_{j,m}lp^{{}^{\prime}}_{ijm}$ $\forall i$ (23)
Additionally, the following set of constraints will ensure that $lp^{{}^{\prime}}_{ijm}=C_{jm}$ if $x_{ijm}=1$ and 00 otherwise in a linear fashion:
$lp^{{}^{\prime}}_{ijm}\leq M^{{}^{\prime}}x_{ijm}$ $\forall i,j,m$ (24)
$lp^{{}^{\prime}}_{ijm}\leq C_{jm}$ $\forall i,j,m$ (25)
$C_{jm}-(1-x_{ijm})M^{{}^{\prime}}\leq lp^{{}^{\prime}}_{ijm}$ $\forall i,j,m$ (26)
Note that $M^{\prime}$ is a big enough positive value.
The objective functions used in this model have not been simultaneously applied to 3D printers in any other research. Moreover, in the literature, the orientation of parts to select their optimal height when allocating them to jobs as shown in this model does not exist with our objectives. Although the orientation of parts in 3D printers has been studied, it has not been addressed in the context of scheduling and production planning following our model.
6 Numerical Experiments
In this section, we solve a numerical example using the epsilon constraint method explained. We expect that as the solution to one of the objective functions improves, the other worsens. To solve this example, we consider two identical 3D printers. There are 9 parts with specified dimensions and delivery deadlines. The details of the parts and machines are provided in the following tables. Also, the example was solved using GAMS studio version 23.4 and CPLEX solver student version 12.8. Table 1 represents the specifications of the machines (printers). Machines are identical.
Table 1: Machine dimensions. Machines are identical in most specifications except for volumetric velocity making these machines nonidentical
Layer Production Time (h/mm) Volumetric Production Time (h/mm3) Dimensions (h×w×l) Machine
0.00006 0.000003 200 × 250 × 250 1
0.00006 0.000003 200 × 250 × 250 2
The dimensions and delivery deadlines of the parts in the numerical example are provided in Table 2 as follows:
Table 2: Dimensions and Delivery Deadlines of Parts in Numerical Example
Part Width (mm) Length (mm) Height (mm) Delivery Deadline (h)
1 5 100 14 24
2 19 19 8 26
3 1.5 70 5 28
4 2 10 9 26
5 10 2.5 2 28
6 3 3.3 8.8 28
7 5 10 10 24
8 3 6 15 24
9 3.5 15 4.5 22
The parts’ dimensions and machine specifications are input into the GAMS software. After solving it with the software, we obtain several solutions. The epsilon constraint method considers the first objective function as the main objective and the other objective function as a constraint with an epsilon limit. The problem is solved by varying the epsilon limits, yielding multiple solutions for different epsilon values.
6.1 Results
Based on the trend shown by Figure 5, it can be understood that the two objective functions conflict although both are minimization problems. As the area increases, the time decreases, and as the time increases, the unused area of the machines decreases. Based on the results, manufacturers can decide on the number of 3D printers needed according to the company’s budget and objectives. They should note that having more 3D printers than necessary increases the related costs, such as maintenance costs, capital holding costs, or lost investment opportunities. Managers should also consider that producing earlier than scheduled may not always benefit their company, as it increases maintenance and capital holding costs. Conversely, having fewer machines than needed can lead to delays in order delivery, resulting in lost customers, decreased popularity, and consequently, lower revenue. Thus, it is based on the company’s policies to decide where to stand on the Pareto front line.
[…]
Figure 5: Pareto Front Diagram
6.2 Scenario Analysis
In this section, we define two scenarios to ensure the accuracy and validity of the presented model. Since the bi-objective model did not yield an overall optimal solution, we removed the second objective function from the model to compare the overall optimal solutions in different situations. In the first scenario, we have constraints for selecting the best height. In the second scenario, we remove these constraints from the model and fix the heights of the parts. We solve these two scenarios for cases with only one machine and two machines with the specifications mentioned in Table 1. For both scenarios, we consider the randomly assigned specifications of parts and due dates as mentioned in Table 3:
Table 3: Dimension specifications of parts and due dates for scenarios
Part Width (mm) Length (mm) Height (mm) Delivery Deadline (h)
1 5 100 140 24
2 19 19 18 26
3 1.5 70 150 28
4 2 10 190 26
5 10 2.5 120 28
6 3 3.3 188 28
7 5 10 110 24
8 3 6 115 24
9 3.5 15 14.5 22
10 25 25 25 24
11 100 100 100 26
12 87 112 112 23
13 90 115 115 43
14 67 116 116 6
15 43 124 124 4
16 32 134 134 25
17 87 144 144 42
18 43 123 123 23
19 43 155 155 21
20 54 142 142 26
Figure 6 represents the sum of earliness and tardiness based on the number of parts for both scenarios. The total earliness and tardiness time in the first scenario is significantly less than in the second scenario. This means that the idea of selecting the height and orientation of parts in the machine when having a single machine with identical parts and production speeds in both scenarios has proven to be very effective. For example, in the first scenario, the sum of earliness and tardiness time for the first $10$ parts is zero, whereas in the second scenario, this value is $17$ hours. The sum of earliness and tardiness time for $20$ parts in the first scenario is $17$ hours, but this value exceeds $70$ hours for the second scenario. Additionally, considering the slope of the curves, as the number of parts increases, the curve of the sum of earliness and tardiness time of parts in the first scenario increases with a lower slope compared to the other scenario. This means that the proposed idea becomes very effective in reducing the sum of earliness and tardiness time for a larger number of parts.
666688881010101012121212141414141616161618181818202020200020202020404040406060606080808080Number of partsSum of earliness and tardinessFirst ScenarioSecond Scenario
Figure 6: Comparison of the sum of earliness and tardiness time between two scenarios when we have one machine
Next, using the data from Table 3, we consider these two scenarios for the case when we have two identical machines. It is expected that in this case, the sum of earliness and tardiness time for both scenarios would decrease as we have more machines. The discussion on which scenario shows a greater reduction in slope will be continued below.
6666888810101010121212121414141416161616181818182020202000101010102020202030303030Number of partsSum of earliness and tardinessFirst ScenarioSecond Scenario
Figure 7: Comparison of the sum of earliness and tardiness between two scenarios when we have two identical machines
According to Figure 7, the sum of earliness and tardiness time has decreased for both scenarios, which was expected based on the number of machines. In this case, as well, the first scenario has performed better than the second scenario. The objective function value in the first scenario was zero for the first $11$ parts, whereas in the second scenario, it was five hours. For $20$ parts, this value is $7$ hours for the first scenario and $29.5$ hours for the second scenario. Also, the increase in the sum of earliness and tardiness for the first scenario when the number of parts is increasing is slower than that of the second scenario.
Next, we will compare the objective function values for both scenarios when we have single-machine and two identical machines.
6666777788889999101010101111111112121212131313131414141415151515161616161717171718181818191919192020202000202020204040404060606060Number of partsSum of earliness and tardinessScenario 1 (Single Machine)Scenario 1 (Two Machines)Scenario 2 (Single Machine)Scenario 2 (Two Machines)
Figure 8: Sum of earliness and tardiness for different scenarios
According to Figure 8, the total delay and lead time for parts in both scenarios, with two machines, have decreased by at least $50$ compared to having only one machine. This reduction is greater for the first scenario than for the second scenario. Specifically, the decrease in the scenario is over $70$ for the first $17$ parts, whereas in the second scenario, this reduction only occurs for the first $10$ parts. An important point to note is that with an increase in the number of parts, the reduction percentage decreases to $50$ for both scenarios. This implies that increasing the number of parts may lead the first scenario also to face machine space constraints, making some desired heights unattainable due to space limitations. However, it is undeniable that the first scenario has performed significantly better than the second scenario. If the number of parts is very high, the model may inevitably choose undesirable heights for parts due to machine space constraints. Nonetheless, the second scenario will not perform better than the first scenario as the second scenario is a feasible solution for the model and sub-optimal generally.
6.3 Sensitivity Analysis
In this part, we focus on the sensitivity analysis of the layering time ($Vt$), which directly affects the selection of parts’ height. Scenarios one and two were defined in the previous section, and here we have excluded the second objective function for scenario comparison again. We have $15$ parts with randomly generated dimension specifications and due dates illustrated by Table 4 and a single machine with the same specifications as before.
Table 4: Dimensions of parts for sensitivity analysis of layering time parameter ($Vt$)
Part Width (mm) Length (mm) Height (mm) Delivery Deadline (h)
1 55 154 170 24
2 160 170 180 17
3 45 80 160 28
4 30 120 190 26
5 100 25 130 28
6 36 43 108 28
7 60 100 130 21
8 30 60 105 35
9 35 140 145 22
10 15 25 125 34
11 90 102 100 36
12 87 122 112 23
13 140 162 115 18
14 34 67 106 16
15 43 46 124 24
Based on Figure 9, the objective function value for both scenarios remains the same for layering times less than $0.001$. This indicates that when the layering speed is very high, the height difference does not significantly affect the tardiness and earliness. However, as the layering speed decreases (layering time exceeds 0.001), the height differences start influencing the objective function value. Scenario one aims to arrange parts in a way that minimizes height differences, leading to a reduction in tardiness and earliness.
$10^{-6}$$10^{-5}$$10^{-4}$$10^{-3}$$10^{-2}$$10^{-1}$00200200200200400400400400Layer production time (h/mm)Sum of earliness and tardiness (h)Scenario 1Scenario 2
Figure 9: Comparison of scenarios based on layer production time. Scenario one aims to arrange parts in a way that minimizes height differences, leading to a reduction in tardiness and earliness.
We allocate the 15 parts with specifications demonstrated by 5 for a single machine under two different machine areas, $A_{m}$. Our goal is to determine the impact of machine capacity on scenarios one and two were defined in the previous section, and here we have excluded the second objective function for scenario comparison again.
Table 5: Dimensions of parts for sensitivity analysis of machine area parameter ($A_{m}$)
Part Width (mm) Length (mm) Height (mm) Delivery Deadline (h)
1 50 150 180 24
2 190 190 180 16
3 15 70 150 28
4 20 100 190 26
5 100 25 120 38
6 30 33 188 28
7 50 100 110 24
8 30 60 115 34
9 35 150 14.5 22
10 25 25 125 14
11 102 100 120 16
12 87 132 112 23
13 190 122 115 13
14 67 34 116 6
15 43 16 124 14
Based on the two Figures 10 and 11 depicting two different machine surface areas for both scenarios, we can draw two conclusions. Both figures demonstrate the better performance of the first scenario provides an optimal solution compared to the scenario. This is indeed natural as the second scenario gives a solution which is a subset of solutions obtained from the first scenario. These results show the validity and effectiveness of the developed model. Also, it means that arranging part heights in a way that minimizes their height differences reduces the total tardiness and earliness. Additionally, the impact of the initial part’s dimension differences is important. In this case, the height differences between parts can be large or small enough that increasing the machine surface area does not necessarily lead to a better solution which is obvious from Figure 11. Thus, managers should invest in 3D printers based on the specifications of parts (similarity and differences in dimensions), i.e., sometimes bigger machines do not necessarily improve the solutions.
222244446666888810101010121212121414141400100100100100200200200200300300300300Number of partsSum of earliness and tardiness (h)Scenario 1Scenario 2
Figure 10: Comparison of scenarios based on number of parts when $A_{m}=4000\ mm^{2}$
222244446666888810101010121212121414141400100100100100200200200200300300300300Number of partsSum of earliness and tardiness (h)Scenario 1Scenario 2
Figure 11: Comparison of scenarios based on number of parts when $A_{m}=9000\ mm^{2}$
Based on Figure 12, it can be observed that increasing the machine surface area will not worsen the optimal solution. This is because the model has sufficient space to choose the appropriate orientation for each part, allowing it to group parts to minimize height differences and reduce their production time. However, part grouping is not solely based on height differences. The model also considers grouping parts with similar delivery deadlines and determines their correct orientation. For instance, when we have 12 parts, the larger area of the machine allows the model to find a better solution but with the addition of the $13^{th}$ part, the increase in the available area of the machine did not affect the solution. This implies that either the dimensions or due date of this part restrict the model from decreasing the earliness and tardiness. Indeed, the addition of the new part is a new restriction that either worsens the solution or does not change it. Another conclusion drawn is that as the number of parts increases, the ratio of machine surface area to the number of parts decreases, limiting the model’s ability to consider all possible scenarios. Therefore, increasing the machine surface area enables the model to explore all part orientation possibilities and choose the best one. For example, for parts number 9, 12, 14, and 15, the scenario where the machine has a larger surface area performs better than the alternative.
22224444666688881010101012121212141414140020202020404040406060606080808080100100100100Number of partsSum of earliness and tardiness (h)Scenario 1, $A_{m}=4000$ $mm^{2}$cenario 1, $A_{m}=9000$ $mm^{2}$
Figure 12: Comparison of scenarios based on number of parts
7 Conclusions
3D printing stands at the forefront of addressing global challenges posed by population growth and resource demands by offering efficient, sustainable production systems. Unlike traditional methods, 3D printing excels in minimizing material usage and environmental impact while enabling customized, on-demand production across industrial, medical, and aerospace sectors. Efficient production planning and scheduling are pivotal in maximizing the advantages of 3D printing. Integrating just-in-time principles with the optimization of part orientation and machine utilization presents a progressive approach to enhancing operational efficiency and cost-effectiveness. Our study contributes to this area by introducing a multi-objective mixed integer linear programming (MOMILP) model designed to optimize 3D printing production schedules, minimizing earliness and tardiness and maximizing 3D printer utilization. This project addresses the complex task of allocating parts with diverse characteristics to jobs and scheduling these jobs across multiple machines. With a focus on minimizing delay and lead time while maximizing machine utilization, the study explores three distinct part placement scenarios. The mathematical model, designed with dual objectives, aims to achieve Pareto optimal solutions due to the inherent trade-offs involved. An innovative aspect of the model involves investigating how selecting a part’s height dimension affects its delivery time, simplifying the analysis to emphasize time-related efficiencies in 3D printing operations. This paper focuses on the allocation of parts with various characteristics to jobs and scheduling these jobs, considering the orientation optimization of parts in machines. Given the dual objectives of the mathematical model presented, a universally optimal solution is not expected. Therefore, a numerical example was provided and solved using the epsilon-constraint method, resulting in the Pareto optimal curve composed of several achievable solutions. One of the ideas proposed in the model is to examine the effect of selecting one dimension of the part as its height on the delivery time. Key findings from our analysis are as follows:
  • 1.
    The proposed approach effectively reduces total part earliness and tardiness time.
  • 2.
    Increasing machine capacity generally improves solution quality, highlighting the critical role of adequate machine resources in achieving optimal outcomes.
  • 3.
    Higher layering speeds mitigate the impact of height differences among parts, whereas slower speeds affect production earliness and tardiness outcomes.
  • 4.
    Increasing the number of printers enhances solutions, although diminishing returns occur when due to part dimensions similarity, discrepancies, and due dates.
Future endeavors could incorporate parts made from various materials, requiring strategic production planning across different machines. Exploring printers with multiple nozzles operating at varying speeds presents new scheduling challenges and opportunities for efficiency gains. Additionally, leveraging sophisticated heuristic methods could optimize solutions for large-sized problems, acknowledging the problem’s NP-hard complexity. Also, integrating cost-related metrics into objective functions would provide insights into overall expense management and sustainability. This project represents a foundational step towards advancing production scheduling in 3D printing, particularly focusing on time efficiency as a pivotal factor for industrial adoption and scalability.
8 Declaration of competing interest
The author affirms that there are no identifiable conflicting financial interests or personal associations that might have seemed to impact the work described in this article.
9 Data and Code availability
Data and code are available upon request.